OpenStudy (anonymous):

Roll a six-sided die 8 times. What is the probability that 6 will appear at least 5 times?

OpenStudy (danjs):

five 6's in only 8 rolls is pretty difficult

OpenStudy (anonymous):

indeed

OpenStudy (danjs):

the probability of rolling a certain number in a single role is 1 out of 6

OpenStudy (danjs):

Say that happens on first role, the probability for the next roll is the same, 1/6 and for all 8 rolls the same

OpenStudy (anonymous):

k

OpenStudy (danjs):

to get 5 of those you need that 1/6 to happen 5 times 1/6 * 1/6 * 1/6 * 1/6 * 1/6

OpenStudy (anonymous):

1/7776

OpenStudy (danjs):

yeah,. that is for rolling all in a row...

OpenStudy (danjs):

the change of not getting 6 is 5/6 for each roll

OpenStudy (anonymous):

ok

OpenStudy (danjs):

so above we considered 5 rolls , want 8 total rolls the 3 left can be misses with a 5/6 chance

OpenStudy (danjs):

so 5 times chance of getting 6 + 3 times chance of missing is getting 'at least' five sixes out of 8 rolls

OpenStudy (danjs):

[1/6 ]^5 * [5/6] ^3

OpenStudy (anonymous):

in other word is 5(9)?

OpenStudy (danjs):

8 rolls total. multiply the probability for each one together. 5 of them should be 1/6 and 8 of them should be 5/6

OpenStudy (anonymous):

i got like 125/1679616

OpenStudy (anonymous):

ok

OpenStudy (danjs):

yeah , small chance

OpenStudy (anonymous):

do that mean i need to add 5,6,7,and 8?

OpenStudy (danjs):

that fraction is about 0.00007

OpenStudy (anonymous):

that number is too long

OpenStudy (anonymous):

perhaps im wrong

OpenStudy (danjs):

were you shown the formula with the n and k , n rolls , k times something occurs

OpenStudy (anonymous):

\[\left(\begin{matrix}n \\ k\end{matrix}\right) p^k(1-p)^{n-k}\] this formula ?

OpenStudy (danjs):

that looks good if you see, it is the same thing we did above

OpenStudy (danjs):

p = success chance, so (1-p) is the chance of fail k = # success n = # failed p=1/6 , (1-p) = 5/6 k= 5 rolls n = 3 rolls

OpenStudy (anonymous):

ok got 0.00461

OpenStudy (kropot72):

That is correct.

OpenStudy (anonymous):

k thx so much

OpenStudy (kropot72):

You're welcome :)